OpenStudy (anonymous):

A radio station had 152 tickets to a concert. They gave away 3 times as many tickets to listeners as to employees. How many tickets did they give away to employees?

OpenStudy (abb0t):

3x + x = 152

OpenStudy (abb0t):

solve for \(\sf \color{Red}{x}\)

OpenStudy (anonymous):

listeners= L employees=E L+E=152 L=3E 3E+E=152 What's the next step in the process?